# Halfway to the Peak: The JWST MIRI 5.6 micron number counts and source population

### v1

@@ 4.2 Number counts
Up to , detection is aided by a negative -correction as the observed 5.6 m probes up the SED towards the stellar 1.6 m bump, which explains the sharper drop in sources at higher redshift. However, we do see a tail out to which corresponds to early massive galaxies.
@@ 4.4 The nature of MIRI sources not in COSMOS2020
As discussed in Section 3.4, while the bulk of our sources (84%) have counterparts in the COSMOS2020 catalog, 108 MIRI are un-matched. We examined these by eye to remove sources that have likely photometric issues such as being driven by a single hot pixel or being residuals of the stripes we saw prior to background subtraction (see Figure 2). After this visual inspection, we are left with 31 reliable MIRI detections not present in the COSMOS2020 catalog. This translates to a source density of 7/sq.arcmin.
To understand why strong MIRI detected sources may not be in the COSMOS2020 catalog consider that the COSMOS2020 source detection is based on a -weighted image (Weaver et al., 2022a). Figure 13 shows the 5 limits of the COSMOS2020 catalog in these six bands, as well as the four IRAC channels. In all cases, we take the limits from the deep parts of COSMOS since these correspond to the locations of our two COSMOS fields. Note that our MIRI F560W depth is comparable to the IRAC channel 1 and 2 depths in this field, while the IRAC channel 3 image is approximately 3 magnitudes shallower (see figure 13). We overlay two potential models that could explain non-inclusion in the COSMOS2020 catalog, but detection in MIRI 5.6 m. One is a star-forming galaxy with (the “Low-mass dusty” model). The other is a star-forming galaxy with (the “High mass dusty” model). Here we use the CSP models from Maraston (2005), assume an age of 300Myr for both galaxies and use the Calzetti law (Calzetti et al., 1994) for dust attenuation. The first model is included because while, previously, low mass galaxies were typically not considered very dusty, JWST has recently found evidence of a population of extremely dusty dwarfs (with masses and up to , Bisigello et al. 2023; see also Pope et al. 2023).
In Figure 13, our model spectra are chosen to both lie below the limits for source detection in COSMOS2020 but also to have MIRI 5.6 m magnitudes on par with those of the sources shown in Figure 14. Both models satisfy these criteria, but they are clearly distinguishable by the IRAC channel 1 brightness. Therefore in Figure 14 we overlay our MIRI images of these 31 sources with the 2 and 3 contours from the COSMOS IRAC channel 1 image. We used the IRAC channel 1 mosaic combining all available COSMOS data, as produced by Annunziatella et al. (2023). We find that 3-7 of the sources have emission in IRAC channel 1. The range given is to distinguish clean detections from ones affected by blending in the IRAC image. The three clean IRAC detections in particular (ID 1-16, 1-107, 1-110) are all consistent with the low-mass dusty model. However, the bulk of our 31 sources are undetected in these IRAC images suggesting they are higher-mass and higher redshift dusty sources. Further study of these 31 sources, folding on the new COSMOS-Web data (Casey et al., 2023), as well as all other available data in COSMOS is reserved for a separate paper.
@@ 5 Summary & Conclusions
In this paper, we present a study of the number counts and source properties of sources detected in eight MIRI F560W images with a combined area of 18.4 sq.arcmin and a 5 depth of 0.1 Jy. Two of our eight fields overlap with the COSMOS2020 catalog which also allows for a more detailed look at the source population properties. Below we summarize our key findings:
The pipeline reduced MIRI F560W images required further background subtraction, given strong stripping artifacts. We also note some residual structure seen in the noise, consistent with tree ring artifacts as noted in documentation. The photometric zero point needed to be adjusted since the header keyword in these early images was not using the true measured MIRI PSF.
Our study includes 8 widely separated fields, allowing us to explore the effects of cosmic variance. We used this to construct field-to-field variation uncertainty on our combined MIRI 5.6 m number counts.
Our number counts have a more pronounced knee, at 2 Jy and are higher than those computed from the CEERS images (Yang et al., 2023). This difference however is only at 1 level given our measured field-to-field variation. The observed counts are consistent with the Cowley et al. (2018) SAM predictions around the knee. These models however overpredict the observed counts below the knee of the counts.
@@
We find 31 very red sources that are securely detected in MIRI, but are not in the COSMOS2020 catalog. This population has a source density of 7/sq.arcmin. They are consistent with being either intermediate redshift () dusty lower mass galaxies (those with a strong IRAC channel 1 detection, 10–20%) or high mass, high- () dusty galaxies (those with weak or no IRAC channel 1 detections, 80–90%). These MIRI 5.6 m only sources will be explored further in a follow-up paper.
@@ Acknowledgments
We are grateful to Meredith Stone for providing us with her MIRI 5.6 m counts ahead of publication. We are also grateful to Matthew Ashby for providing us with a copy of his old IRAC channel 3 counts. Based on observations with the NASA/ESA/CSA JWST obtained at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Incorporated, under NASA contract NAS5-03127. Support for program number JWST-GO-01762 was provided through a grant from the STScI under grant number JWST-GO-01762.010-A.

### v2

@@ 4.3 The MIRI sources in the COSMOS2020 catalog
Up to , detection is aided by a negative -correction as the observed probes up the SED towards the stellar bump, which explains the sharper drop in sources at higher redshift. However, we do see a tail out to which corresponds to early massive galaxies.
@@ 4.4 The nature of MIRI sources not in COSMOS2020
As discussed in Section 3.4, while the bulk of our sources (84%) have counterparts in the COSMOS2020 catalog, 108 MIRI sources are un-matched. We examined these by eye to remove sources that have likely photometric issues such as being driven by a single hot pixel or being residuals of the stripes we saw prior to background subtraction (see Figure 2). After this visual inspection, we are left with 31 reliable MIRI detections not present in the COSMOS2020 catalog. This translates to a source density of 7 per square arcminute.
To understand why strong MIRI detected sources may not be in the COSMOS2020 catalog consider that the COSMOS2020 source detection is based on a -weighted image (Weaver et al., 2022a). Figure 14 shows the 5 limits of the COSMOS2020 catalog in these six bands, as well as the four IRAC channels. In all cases, we take the limits from the deep parts of COSMOS since these correspond to the locations of our two COSMOS fields. Note that our MIRI F560W depth is comparable to the IRAC channel 1 and 2 depths in this field, while the IRAC channel 3 image is approximately 3 magnitudes shallower (see figure 14). We overlay two potential models that could explain non-inclusion in the COSMOS2020 catalog, but detection in MIRI . One is a star-forming galaxy with (the “Low-mass dusty” model). The other is a star-forming galaxy with (the “High mass dusty” model). Here we use the Composite Stellar Population (CSP) models from Maraston (2005), assume an age of 300Myr for both galaxies and use the Calzetti law (Calzetti et al., 1994) for dust attenuation. The first model is included because while, previously, low mass galaxies were typically not considered very dusty, JWST has recently found evidence of a population of extremely dusty dwarfs (with masses and up to , Bisigello et al. 2023; see also Pope et al. 2023).
In Figure 14, our model spectra are chosen to both lie below the limits for source detection in COSMOS2020 but also to have MIRI magnitudes on par with those of the sources shown in Figure 15. Both models satisfy these criteria, but they are clearly distinguishable by the IRAC channel 1 brightness. Therefore in Figure 15 we overlay our MIRI images of these 31 sources with the 2 and 3 contours from the COSMOS IRAC channel 1 image. We used the IRAC channel 1 mosaic combining all available COSMOS data, as produced by Annunziatella et al. (2023). We find that 3-7 of the sources have emission in IRAC channel 1. The range given is to distinguish clean detections from ones affected by blending in the IRAC image. The three clean IRAC detections in particular (ID 1-16, 1-107, 1-110) are all consistent with the low-mass dusty model. However, the bulk of our 31 sources are undetected in these IRAC images suggesting they are higher-mass and higher redshift dusty sources. Further study of these 31 sources, folding on the new COSMOS-Web data (Casey et al., 2023), as well as all other available data in COSMOS is reserved for a separate paper.
@@ 5 Summary & Conclusions
In this paper, we present a study of the number counts and source properties of sources detected in eight MIRI F560W images with a combined area of 18.4 sq.arcmin and a 5 depth of . Two of our eight fields overlap with the COSMOS2020 catalog which also allows for a more detailed look at the source population properties. Below we summarize our key findings:
The pipeline reduced MIRI F560W images required further background subtraction, given strong striping artifacts. We also note some residual structure seen in the noise, consistent with tree ring artifacts as noted in documentation. The photometric zero point needed to be adjusted since the header keyword in these early images was not using the true measured MIRI PSF.
Our study includes 8 widely separated fields, allowing us to explore the effects of cosmic variance. We used this to construct field-to-field variation uncertainty on our combined MIRI number counts.
Our number counts have a more pronounced knee, at and are higher than those computed from the CEERS images (Yang et al., 2023). This difference however is only at 1 level given our measured field-to-field variation. The observed counts are consistent with the Cowley et al. (2018) SAM predictions around the knee. These models however overpredict the observed counts below the knee of the counts.
@@
We find 31 very red sources that are reliably detected in MIRI, but are not in the COSMOS2020 catalog. This population has a source density of 7 per square arcminute. They are consistent with being either intermediate redshift () dusty lower mass galaxies (those with a strong IRAC channel 1 detection, 10 to 20%) or high mass, high- () dusty galaxies (those with weak or no IRAC channel 1 detections, 80 to 90%). These MIRI only sources will be explored further in a follow-up paper.
@@ Acknowledgments
We are grateful to the anonymous referee for their careful reading of our manuscript and detailed feedback which greatly improved the clarity of this paper. We are grateful to Meredith Stone for providing us with her MIRI 5.6 m counts ahead of publication. We are also grateful to Matthew Ashby for providing us with a copy of his old IRAC channel 3 counts. Based on observations with the NASA/ESA/CSA JWST obtained at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Incorporated, under NASA contract NAS5-03127. Support for program number JWST-GO-01762 was provided through a grant from the STScI under grant number JWST-GO-01762.010-A.